Demand curve: P = 30 – Q Supply curve: P = 2Q

Calculate the equilibrium quantity and price.
Answer #1

At equilibrium,

Demand = Supply

30 - Q = 2Q

3Q = 30

Q = 30 / 3 = 10

P = 2Q = 2 * 10 = $20

Thus, thev equilibrium quantity is 10 and price is $20.